# Controlling Skyrmion Lattice Orientation with Local Magnetic Field Gradients

### v1

@@ Abstract
Duc Minh Tran: Conceptualization (equal); Data curation (lead); Investigation (lead); Formal analysis (lead); Methodology (lead); Visualization (equal); Writing – original draft (lead); Writing – review & editing (equal). Elizabeth M. Jefremovas: Conceptualization (equal); Supervision (equal); Data curation (equal); Writing – review & editing (equal). Edoardo Mangini: Investigation (equal); Formal analysis (equal); Methodology (equal). Fabian Kammerbauer: Methodology (equal). Dennis Meier: Methodology (equal); Writing – review & editing (equal). Robert Frömter: Supervision (equal); Visualization (equal); Writing – review & editing (equal). Mathias Kläui: Conceptualization (lead); Resources (lead); Supervision (lead); Writing – review & editing (equal).
@@
The data that support the findings of this study are available from the corresponding author upon reasonable request.

### v2

@@ Abstract
Duc Minh Tran: Conceptualization (equal); Data curation (lead); Investigation (lead); Formal analysis (lead); Methodology (lead); Visualization (equal); Writing – original draft (lead); Writing – review & editing (equal). Edoardo Mangini: Investigation (equal); Formal analysis (equal); Methodology (equal). Elizabeth M. Jefremovas: Conceptualization (equal); Supervision (equal); Data curation (equal); Writing – review & editing (equal). Fabian Kammerbauer: Methodology (equal); Writing – review & editing (equal). Dennis Meier: Methodology (equal); Writing – review & editing (equal). Robert Frömter: Supervision (equal); Visualization (equal); Writing – review & editing (equal). Mathias Kläui: Conceptualization (lead); Resources (lead); Supervision (lead); Writing – review & editing (equal).
@@
The data supporting the findings in this study is publicly available via Zenodo at https://doi.org/10.5281/zenodo.17063771.